If alfa, beta are the zeroes of (x-l)(x-m)-n, prove that l,m are zeroes of (x-alpha)(x-beta)+n
 

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
(x - alpha) (x - beta) = (x-l) (x-m) -n
 
Now
 
(x- alpha) (x- beta) +n = (x-l) (x-m)
 
Hence l and m are roots